Using traditional methods, it takes 11.2 hours to receive a basic flying license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 16 students and observed that they had a mean of 11.4 hours with a standard deviation of 1.6. A level of significance of 0.10 will be used to determine if the technique performs differently than the traditional method. Assume the population distribution is approximately normal. Find the value of the test statistic. Round your answer to three decimal places.
Answer #1

Solution :

\mu= 11.2

\bar x = 11.4

s = 1.6

n = 16

\alpha = 0.10

Test statistic = t =

= (\bar x - \mu ) / s / \sqrt n

= (11.4 -11.2 ) / 1.6/ \sqrt 16

Test statistic = t = 0.500